OpenStudy (anonymous):

Integral of e^(5t+pi)

OpenStudy (raden):

ntegral by sub.. let u=5t+pi du = 5 dt or dt = du/5 so, the integration can be : int e^(5t+pi) dt (i assumed it respect to dt) =1/5 * int e^u du = ...

OpenStudy (raden):

just remember the formula int e^x = e^x too

OpenStudy (anonymous):

=1/5*int t e^u du ... and then I used Integration by Parts? I got (1/5 e^(5t+pi))(t-1) + c Is this right?

OpenStudy (raden):

here, we neednt use int by parts, but it is integration by substitution ...

OpenStudy (anonymous):

Yes, I used int by sub first. Then, I used parts.

OpenStudy (raden):

1/5 * int e^u du = 1/5 * e^u + c subt back that u=5t+pi , therefore = 1/5*(e^5t+pi ) + c

OpenStudy (raden):

hmm..typo = 1/5*e^(5t+pi ) + c

OpenStudy (anonymous):

Sorry! I didn't give you the rest of the problem. That's why I am confused. Integral of t * e^(5t + pi) dt

OpenStudy (anonymous):

=1/5*int t e^u du ... and then I used Integration by Parts? I got (1/5 e^(5t+pi))(t-1) + c Thank you so much for helping me.

OpenStudy (raden):

Oooooo ....~0.0~

OpenStudy (anonymous):

Sorry about that. I asked you for just the e^ part and then the problem had an extra (t) in it!! Thank you for helping me!!

OpenStudy (raden):

nevermind... im a true helper :) well, we need int by part